A-   For all integers a and b, if a + b is odd, then a is odd or b is odd.

B-  For any integer n the number (n3 - n) is even.

C-   Proof of De-Morgan’s Law


1
Expert's answer
2021-11-08T16:52:17-0500

A. For all integers "a"  and "b" let "a+b" is odd. If "a" and "b" are both even than "\exist p, l \in Z a=2p, b=2l\implies a+b=2(p+l)" is even too. If "a" and "b" are both odd than "\exist p, l \in Z a=2p+1, b=2l+1 \implies a+b=2(p+l+1)" is even too. So we have only one of  "a"  and "b" is odd and "\exist p, l \in Z a=2p, b=2l+1 \implies a+b=2(p+l)+1" .

B. For any integer "n"  the number "(n^3-n)" equals to "n(n+1)(n-1)" that is the product of three three consecutive natural numbers, one of which is even. In case "(n-1)=0" we have "(n^3-n)=0" and it is even too.

C. Let's prove that for alll sets "A" and "B" the equality "\lnot(A\land B)=\lnot A\lor \lnot B" .

"\lnot (A\land B)=(\lnot A\land B)\lor ( A\land \lnot B) \lor ( \lnot A\land \lnot B) ="

"((\lnot A\land B) \lor ( \lnot A\land \lnot B))\lor (( A\land \lnot B) \lor ( \lnot A\land \lnot B)) ="

"( \lnot A \land ( B\lor \lnot B))\lor (\lnot B \land ( B\lor \lnot B)) =\lnot A\lor \lnot B"